A transverse wave on a string is modeled with the wave function y(x, t) (0.80 m)sin[(0.85 m)x (1.70 s)t 0.20]. (Indicate the direction with the signs of your answers.) (a) Find the wave velocity (in m/s). m/s (b) Find the position (in cm) in the y-direction, the velocity (in cm/s) perpendicular to the motior of the wave, and the acceleration (in cm/s2) perpendicular to the motion of the wave of a small segment of the string centered at x 0.40 m at time t- 5.00 s position velocity acceleration cm cm/s cm/s2
Answer #1

Given wave equation for displacement is y(x, t) = (0.80) sin(0.85x + 1.70t 0.20)

k 0,85 m-1 , 1.70 s-1

a)

Wave velocity w1.70 u=-=0.85 = 2 m/s

b)

y(x, t) = (0.80) sin(0.85x + 1.70t 0.20)

At 0.40 771 and t 5.00 s

Position  y(0.80)sin(0.85(0.40)1.70 5.00) + 0.20) 0.30 m 30 cm

Velocity u(z, t) = y(z, t) = (0.80) (1.70) cos(0.85x + 1.70t + 0.20)

At 0.40 771 and t 5.00 s

v (0.80) ( 1.70) cos(0.85(0.40)+1.70(5.00)+0.20)--1.26 m/s--126 ст/s

Acceleration a(x, t) =-u(r, t) =-(0.80)(1.70)2 sin(0.85x + 1.70t + 0.20) dt

At 0.40 771 and t 5.00 s

a =-(0.80)( 1.70)-sin( 0.85(0.40) + 1.70( 5.00) +0.20) =ー0.868 m/s =- 86.8 cm s